OpenStudy (anonymous):

Help! I need to show work for this question. 12b^-2-7b^-1+1=0

OpenStudy (anonymous):

I got b = 4,3

OpenStudy (anonymous):

\[\frac{ 12 }{ b^{2} }-\frac{ 7 }{ b }+1=0\] Multiply both sides by b^2 \[b^{2}-7b+12=0\] So your answer looks right.